# neurips_2025.bib

### v1

@@ 4 Experiment
4.1 Main Results for Post-training
Table 1: Performance comparison on Google Robot and WidowX Robot in SimplerEnv [simplerenv]. The experiments are conducted across 12 tasks, including both visual matching (VM) and visual aggregation (VA) settings. The “(x B)” following the method name indicates the parameter size of the large language model. The percentage sign % of the success rate is omitted.
@@ 4.1 Main Results for Post-training
RT-1-X [oxe] 59.7 29.4 21.3 10.1 56.7 49.0 31.7 32.3 42.4 30.2 0.0 4.2 0.0 0.0 1.1 24.6
RT-2-X [oxe] 25.0 35.5 3.7 20.6 78.7 82.3 77.9 79.2 46.3 54.4 - - - - - -
Octo-Base [octo] 22.7 1.1 0.0 0.0 17.0 0.6 4.2 3.1 11.0 1.2 15.8 12.5 0.0 41.7 17.5 9.9
RoboVLMs (2B) [robovlms] 44.9 10.3 27.8 0.0 76.3 50.7 79.0 62.5 57.0 30.9 41.7 33.3 0.0 62.5 34.4 40.8
SpatialVLA (3B) [spatialvla] 54.6 39.2 0.0 6.3 79.3 78.7 90.0 83.0 56.0 51.8 20.8 37.5 41.7 83.3 45.8 51.2
TraceVLA-Phi3 (4B) [Tracevla] 35.4 37.5 0.0 0.0 69.7 75.4 70.8 67.8 44.0 45.1 8.3 0.0 12.5 66.7 21.9 37.0
CronusVLA (0.5B) 50.5 36.8 42.6 21.7 96.0 94.6 93.0 78.0 70.5 57.8 45.8 33.3 0.0 79.2 39.6 56.0
OpenVLA (7B) [openvla] 59.7 23.5 0.0 2.9 25.7 54.1 55.0 63.0 35.1 35.9 8.3 4.2 0.0 0.0 3.1 24.7
CogACT (7B) [cogact] 71.8 28.3 50.9 46.6 91.3 89.6 85.0 80.8 74.8 61.3 71.7 50.8 15.0 67.5 51.3 62.4
TraceVLA (7B) [Tracevla] 63.1 61.6 11.1 12.5 45.0 64.3 63.8 60.6 45.8 49.8 12.5 16.6 16.6 65.0 27.7 41.1
Magma (8B) [magma] 58.9 59.0 8.3 24.0 75.0 68.6 53.0 78.5 48.8 57.5 37.5 29.2 20.8 91.7 44.8 50.4
Basic-Post (7B) [openvla] 66.7 36.0 2.8 7.4 34.3 53.9 69.2 60.4 43.3 39.4 8.3 0.0 12.5 20.8 10.4 31.0
CronusVLA (7B) 77.8 58.7 64.8 65.1 95.7 94.2 76.0 77.0 78.6 73.8 66.7 54.2 20.8 100.0 60.4 70.9
@@
Implementation details. Our focus is on exploring multi-frame modeling during post-training, building on standard off-the-shelf pretraining methods. In this section, we primarily investigate the performance of our post-trained model. After pretraining the basic single-frame VLA following [openvla, belkhale2024minivla] with the OXE dataset [oxe], we select two high-quality datasets, Bridge-v2 [bridgev2] and Fractal [RT1] datasets, to conduct further cross-embodiment post-training with multi-frame modeling, which include about 148k episodes and 5M multi-frame clips. Our CronusVLA 7B is built on 7B Llama 2 [llama2], and CronusVLA 0.5B is built on Qwen2.5 0.5B [yang2024qwen2]. Following [prismatic], they both employ the Dinov2  [dinov2] and SigLip [siglip] as vision encoders. For input condition, CronusVLA is built on a third-person camera and a text instruction. In addition to the current one-frame observation, CronusVLA-7B is configured with a default of 6 past frames, while CronusVLA-0.5B uses 3 past frames. All experiments are based on A100 GPUs. More training details are included in Appendix C.
SimplerEnv and baselines. We conduct simulation experiments within SimplerEnv [simplerenv], a benchmark designed to evaluate the models in performing various tasks with the WidowX Robot (WR) and the Google Robot (GR) environment. The GR environment includes two experimental settings, Visual Matching (VM) and Variant Aggregation (VA), one strictly follows a real-to-sim replication, and another introduces environmental variations. WidowX Robot environment only includes the VM setting. Visualizations are shown in Figure 4. We report the average success rate and adopt the same evaluation setup as in [cogact]. For pretrained models, RT-1-X [oxe, RT1], RT-2-X [oxe, RT2], and Octo-Based [octo] are early baselines, OpenVLA [openvla], CogACT [cogact], and Magma [magma] are trained on subsets of the OXE dataset and have model sizes exceeding 7B parameters. For the remaining reported models, they are all post-trained on the Fractal and Bridge-v2 datasets. RoboVLMs (2B)[robovlms] is a multi-frame VLA model; we report the results of two official checkpoints trained separately on the two datasets, evaluated independently on the Google and WidowX robots. SpatialVLA (official mixture version)[spatialvla] is post-trained from a pretrained version. TraceVLA (7B) and TraceVLA-Phi3 (4B) [Tracevla] are trained with additional visual prompting annotations. Basic-Post (7B) denotes the model post-trained in a discrete manner, which starts from our pretrained basic model [openvla]. All models are evaluated using official checkpoints, or their official results are directly reported.
@@
Results in SimplerEnv. Main results are illustrated in Table 1. For Google Robot setting, our CronusVLA-7B achieves the highest average success rate, with 78.6 in the VM and 73.8 in the VA, surpassing TraceVLA and RoboVLMs, also multi-frame VLAs, by +71.6% and +37.9% relative VM score, +48.2% and +138.8% relative VA scores. Notably, our model achieves strong performance on not only simple tasks Pick Coke Can and Open/Close Drawer, but also the more complex task Put in Drawer, a long-horizon task, requires sequential actions of opening the drawer and placing the apple in. Most previous approaches have consistently failed to attain high success rates. Our method effectively improves the VM success rate to 64.8 and the VA to 65.1 in this task. The highest average scores of VA and VM show our better performance and robustness. For the WidowX robot, CronusVLA 7B shows superior performance and achieves the highest average success rate, +41.5% higher than SpatialVLA and +17.7% higher than CogACT. CronusVLA 0.5B, with a 0.5B language model, outperforms many prior models trained on larger (from 2B to 7B) language models. It achieves the best results in Pick Coke Can and Move Near tasks over all other models, suggesting that excessive parameters may not always be beneficial for simple tasks, emphasizing the value of effective modeling. Both CronusVLA 0.5B and 7B show a promising performance in SimplerEnv.
Figure 4: Visualization of SimplerEnv [simplerenv] and LIBERO [Libero] simulation settings.
Table 2: Main results in LIBERO [Libero], the average success rate across 3 seeds over 500 trials per task.
LIBERO Spatial Object Goal Long Ave.
Diffusion Policy [diffusionpolicy] 78.3 92.5 68.3 50.5 72.4
MDT [reuss2024multimodal] 78.5 87.5 73.5 64.8 76.1
Octo [octo] 78.9 85.7 84.6 51.1 75.1
OpenVLA (7B) [openvla] 84.7 88.4 79.2 53.7 76.5
TraceVLA (7B) [Tracevla] 84.6 85.2 75.1 54.1 74.8
SpatialVLA (3B) [spatialvla] 88.2 89.9 78.6 55.5 78.1
CronusVLA (0.5B) 90.4 89.8 83.2 59.5 80.7
CronusVLA (7B) 90.1 94.7 91.3 68.7 86.2
4.2 Main Results for Finetuning
Evaluation Setup in LIBERO. We evaluate our finetuning stage in the LIBERO [Libero] simulation benchmark. LIBERO comprises four task suites, including LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long, and they separately evaluate spatial reasoning, object and goal generalization, and long-horizon planning from precise placement to multi-step goal execution. Based on our multi-frame post-trained models, we conduct finetuning for each suite with the action adaptation strategy. We compare our method with the low-level policy Diffusion Policy  [diffusionpolicy], MDT  [reuss2024multimodal], both of which are trained from scratch. Other models that are finetuned from pre-trained weights, including Octo [octo], OpenVLA [openvla], TraceVLA [Tracevla], SpatialVLA [spatialvla]. All above models except MDT are conditioned on one third-person observation and language instruction, without multi-view observations and state. CronusVLA 7B is finetuned 10K steps starting from our post-trained weight, and the other 5K~12K steps are employed to train the action adapter. The past frame number is 3.
Results in LIBERO. Figure 4 presents the evaluation results in the LIBERO benchmark. Our CronusVLA (7B) achieves the highest average success rate of 86.2%, surpassing existing baselines, including strong generalist models such as SpatialVLA (78.1%), OpenVLA (76.5%), and TraceVLA (74.8%). Notably, CronusVLA 7B attains good performance across almost all individual task suites, achieving 94.7% on Object, 91.3% on Goal, and a remarkable 68.7% on long-horizon tasks—demonstrating its strong capability in learning long-horizon executions under multi-frame modeling and action adaptation. In addition, CronusVLA 0.5B, despite its substantially smaller backbone, outperforms several larger models with an average score of 80.7%. It sets the best performance on LIBERO-Spatial (90.4%), showcasing the effectiveness of our model design even under limited capacity. These results demonstrate the effectiveness of our multi-frame architecture and training strategy for downstream task finetuning.
@@ 4.2 Main Results for Finetuning
Figure 5: Real-world experiment with Franka platform. Evaluation of basic pick-and-place capabilities is in (a), long-horizon tasks in (b) demonstrate the advantages of multi-frame modeling in handling temporally dependent manipulations, and (c) generalization and robustness tests, particularly under camera occlusion and various disturbances, highlight the robustness of our model.
@@
Evaluation setup with Franka platform. As shown in Figure 5, we evaluate our method on several real-world tasks with the Franka Research 3 Robot, and utilize the delta end-effector action to control Franka; meanwhile, we utilize a third-person camera for visual input. Three task suites are designed: (1) Simple pick-and-place, involves the picking and placing objects with varying colors and shapes across different locations and orientations; (2) Long-horizon tasks, requires coordinated multi-step manipulation and includes putting multiple objects, opening the drawer then closing it, placing objects into a drawer and pressing buttons in a specific order; and (3) Generalization and robustness tasks, evaluating performance on unseen objects, novel instructions, camera occlusions, distractor objects and so on. We manually collect 30 demonstration episodes for each pick objects task, and 50 episodes for the other tasks. All expert demonstrations are used for co-training, and the number of successful rollouts of 25 trials is reported. We implement 3D Diffusion Policy (DP3) [dp3], and OpenVLA [openvla] finetuned on these demonstrations, our CronusVLA 7B is finetuned from the post-trained weight. More details of deployment and task descriptions are in Appendix E.
@@
Results with Franka platform. CronusVLA outperforms other models across almost all tasks. For simple pick-and-place tasks in Figure 5 (a), all three policies perform well when handling simple objects; however, for tasks requiring more precise manipulation, such as grasping the edge of a bowl or stacking one block (or cup) on another, CronusVLA shows better in-domain performance. Long-horizon tasks are shown in Figure 5 (b). CronusVLA exhibits stronger long-horizon learning capabilities, achieving consistently better performance than DP3 and OpenVLA under limited expert demonstrations. Notably, in press buttons in order, OpenVLA tends to press the same button multiple times, indicating state confusion in long-horizon tasks due to the absence of multi-frame information. CronusVLA has inherent temporal awareness to effectively handle state-ambiguous scenarios for button pressing, and has better perception ability for the positioning and orientation required for opening or closing drawers. Generalization and robustness tasks are illustrated in Figure 5 (c) and (d). In all designed distracted situations, our model shows more general and robust capability than DP3 and OpenVLA. DP3 is sensitive to distractions from extraneous objects and human interference. In the Camera Occlusions task, both DP3 and OpenVLA are adversely affected by frequent visual input dropouts, leading to performance degradation due to their reliance on precise observations for every step. In contrast, our multi-frame modeling effectively withstands such disturbances. We attribute this robustness to the adaptive selection capability of our cross-attention mechanism in the cross-frame decoder, which can prevent corrupted current observations from inducing out-of-distribution actions and also can mitigate the influence of noisy past observations on current execution.

### v2

@@ 4 Experiment
4.1 Results in Simulation
@@
RT-1-X (O’Neill et al. 2024) 59.7 29.4 21.3 10.1 56.7 49.0 31.7 32.3 42.4 30.2 0.0 4.2 0.0 0.0 1.1 24.6
RT-2-X (O’Neill et al. 2024) 25.0 35.5 3.7 20.6 78.7 82.3 77.9 79.2 46.3 54.4 - - - - - -
Octo-Base (Octo Model Team et al. 2024) 22.7 1.1 0.0 0.0 17.0 0.6 4.2 3.1 11.0 1.2 15.8 12.5 0.0 41.7 17.5 9.9
RoboVLMs (2B) (Li et al. 2024c) 44.9 10.3 27.8 0.0 76.3 50.7 79.0 62.5 57.0 30.9 50 37.5 0.0 83.3 42.7 43.5
SpatialVLA (3B) (Qu et al. 2025) 54.6 39.2 0.0 6.3 79.3 78.7 90.0 83.0 56.0 51.8 20.8 37.5 41.7 83.3 45.8 51.2
$\pi_{0}$ (3B) (Black et al. 2024) 38.3 25.6 - - 72.7 75.2 65.3 63.7 - - 25.0 16.7 12.5 29.2 20.9 -
$\pi_{0}$-FAST (3B) (Pertsch et al. 2025) 42.9 31.3 - - 75.3 77.6 67.5 68.2 - - 29.1 21.9 10.8 66.6 32.1 -
GR00T-N1.5 (2B) (Bjorck et al. 2025) 27.8 35.8 7.4 4.0 51.7 69.3 54.0 68.7 35.2 44.5 75.3 54.3 57.0 61.3 61.9 47.2
[…]
OpenVLA (7B) (Kim et al. 2024) 59.7 23.5 0.0 2.9 25.7 54.1 55.0 63.0 35.1 35.9 8.3 4.2 0.0 0.0 3.1 24.7
CogACT (7B) (Li et al. 2024a) 71.8 28.3 50.9 46.6 91.3 89.6 85.0 80.8 74.8 61.3 75.0 50.0 16.7 79.2 55.2 63.8
TraceVLA (7B) (Zheng et al. 2024) 63.1 61.6 11.1 12.5 45.0 64.3 63.8 60.6 45.8 49.8 12.5 16.6 16.6 65.0 27.7 41.1
OpenVLA-OFT (7B) (Kim, Finn, and Liang 2025) 47.2 12.2 0.9 0.5 72.3 65.3 69.6 59.0 47.5 34.3 12.5 4.2 4.2 100.0 30.2 37.3
Magma (8B) (Yang et al. 2025a) 58.9 59.0 8.3 24.0 75.0 68.6 53.0 78.5 48.8 57.5 37.5 29.2 20.8 91.7 44.8 50.4
[…]
Table 1: Performance comparison on SimplerEnv. Experiments are conducted across 12 tasks. Sizes of the original LLMs are listed, with % omitted. indicates evaluating all 12 tasks using a co-trained checkpoint; otherwise, two checkpoints are used to evaluate the Google Robot and WidowX setups separately.
@@
Implementation details. In this section, we primarily investigate the performance of our post-trained model. After pretraining the basic single-frame VLA following (Kim et al. 2024; Belkhale and Sadigh 2024) with the OXE dataset (O’Neill et al. 2024), we select two high-quality datasets, Bridge-v2 (Walke et al. 2023) and Fractal (Brohan et al. 2022) datasets, to conduct multi-frame post-training, which include about 148k episodes and 5M multi-frame clips. Our CronusVLA 7B is built on 7B Llama 2 (Touvron et al. 2023), and CronusVLA 0.5B is built on Qwen2.5 0.5B (Yang et al. 2024a), they both employ Dinov2 and SigLip as vision encoders. CronusVLA is built on a third-person camera and a text instruction. CronusVLA 7B is configured with a default of 6 past frames, while CronusVLA 0.5B utilizes 3 past frames. All experiments are based on A100 GPUs. More training details are included in Appendix.
Evaluation setup in SimplerEnv. We conduct experiments within SimplerEnv (Li et al. 2024b), a benchmark to evaluate models across various tasks with the WidowX Robot (WR) and the Google Robot (GR). The GR environment includes two settings, Visual Matching (VM) and Variant Aggregation (VA). WidowX Robot environment only includes the VM setting. SimplerEnv covers more than 2k trails across different scenarios, objects, and tasks. We report the average success rate of each task. RT-1-X (O’Neill et al. 2024; Brohan et al. 2022), RT-2-X (O’Neill et al. 2024; Brohan et al. 2023), and Octo-Based (Octo Model Team et al. 2024) are early baselines, OpenVLA (Kim et al. 2024), CogACT (Li et al. 2024a), and Magma (Yang et al. 2025a) are trained on subsets of the OXE. $\pi_{0}$, $\pi_{0}$-FAST and GR00T-N1.5 are VLAs with additional robot state input. RoboVLMs (Li et al. 2024c), SpatialVLA (Qu et al. 2025), and TraceVLA (Zheng et al. 2024) are all trained on the Fractal and Bridge-V2 datasets.
@@
Experimental results on SimplerEnv. Main results are illustrated in Tab.1. For Google Robot setting, our CronusVLA-7B achieves the highest average success rate, with 78.6 in the VM and 73.8 in the VA, surpassing TraceVLA and RoboVLMs (also multi-frame VLAs), by +71.6% and +37.9% relative VM score, +48.2% and +138.8% relative VA scores. Notably, our model achieves strong performance on not only simple tasks Pick Coke Can and Open/Close Drawer, but also the long-horizon task Put in Drawer, which requires sequential actions of opening the drawer and placing the apple in it. Most previous approaches have failed to attain high success rates in Put in Drawer, while our method effectively improves its VM success rate to 64.8 and the VA to 65.1. For the WidowX, CronusVLA 7B also achieves a high average success rate, +41.5% higher than SpatialVLA and +9.4% higher than CogACT. CronusVLA 0.5B, with a smaller language backbone, outperforms many prior models trained on larger (from 2B to 7B) language models. It achieves the best results in Pick Coke Can and Move Near over all other models, suggesting that excessive parameters may not always be beneficial, emphasizing the value of effective modeling.
Methods Spatial Object Goal Long Ave.
Diffusion Policy (Chi et al. 2023) 78.3 92.5 68.3 50.5 72.4
Dita (Hou et al. 2025) 84.2 96.3 85.4 63.8 82.4
OpenVLA (Kim et al. 2024) 84.7 88.4 79.2 53.7 76.5
TraceVLA (Zheng et al. 2024) 84.6 85.2 75.1 54.1 74.8
SpatialVLA (Qu et al. 2025) 88.2 89.9 78.6 55.5 78.1
$\pi_{0}$-FAST (Pertsch et al. 2025) 96.4 96.8 88.6 60.2 85.5
GR00T-N1 (Bjorck et al. 2025) 94.4 97.6 93.0 90.6 93.9
$\pi_{0}$ (Black et al. 2024) 96.8 98.8 95.8 85.2 94.2
$\pi_{0.5}+KI$ (Intelligence et al. 2025) 98.0 97.8 95.6 85.8 94.3
[…]
Table 2: Main results in LIBERO (Liu et al. 2024), the average success rates are reported. More details are in Tab.16.
Temporal Dimension WidowX Robot
Constant (1:0) Cyclic (1:1) Sparse (1:3) VM
Methods R-Score SR R-Score SR R-Score SR - SR
$\pi_{0}$ (Black et al. 2024) 43.5 9.1 36.8 7.7 34.9 7.3 - 20.9
TraceVLA (Zheng et al. 2024) 59.2 16.4 62.5 17.3 78.0 21.6 - 27.7
RoboVLMs (Li et al. 2024c) 47.6 20.3 57.3 24.5 78.7 33.6 - 42.7
SpatialVLA (Qu et al. 2025) 44.9 20.6 48.0 22.0 63.1 28.9 - 45.8
CogACT (Li et al. 2024a) 53.3 29.4 66.1 36.5 80.2 44.3 - 55.2
[…]
Spatial Dimension Total Avg.
Global Local Discrete
Methods R-Score SR R-Score SR R-Score SR R-Score SR
$\pi_{0}$  (Black et al. 2024) 42.6 8.9 28.2 5.9 52.1 10.9 41.1 8.6
TraceVLA (Zheng et al. 2024) 58.1 16.1 65.3 18.1 81.9 22.7 65.8 18.2
RoboVLMs (Li et al. 2024c) 54.7 23.4 83.3 35.6 76.8 32.8 67.4 28.8
SpatialVLA (Qu et al. 2025) 57.6 26.4 50.0 22.9 52.4 24.0 54.4 24.9
CogACT (Li et al. 2024a) 60.2 33.2 80.5 44.4 87.4 48.3 72.1 39.8
[…]
Table 3: Observational robustness test on SimplerEnv-OR. Top: Temporal dimension with different frequencies. Bottom: Spatial dimension with different patterns. R-Score is the pre-defined score of robustness. SR denotes the success rate (%).
@@
Figure 4: Real-world experiment. Evaluation of basic pick-and-place capabilities is in (a), long-horizon tasks of (b) demonstrate the advantages of multi-frame modeling in handling temporally dependent manipulations, and (c) generalization and robustness tests, particularly under camera occlusion and various disturbances, highlight the robustness of our model.
Evaluation setup and experimental results in LIBERO. We evaluate CronusVLA on the LIBERO (Liu et al. 2024) benchmark. LIBERO comprises four task suites, including Spatial, Object, Goal, and Long. Based on the post-trained weights, CronusVLA is fully finetuned on each suite. We compare it with Diffusion Policy (Chi et al. 2023), Dita (Hou et al. 2025), OpenVLA, TraceVLA, and SpatialVLA, as well as other models conditioned on an additional wrist-view image and the robot state, including $\pi_{0}$, $\pi_{0}$-FAST, GR00T-N1, and $\pi_{0.5}$. As shown in Tab.2, our model with only an extra wrist-view input achieves the highest average success rate of 97.0%, attaining good performance across almost all individual suites and a remarkable 94.0% on Long (a +40.3% over OpenVLA), which demonstrates our robust in-domain learning and long-horizon execution capabilities.
Robustness testing on SimplerEnv-OR. We evaluate the observational robustness of several models: RoboVLMs (also a multi-frame VLA), SpatialVLA (spatial modeling VLA), CogACT (prior SOTA model), TraceVLA (designed for temporal modeling), $\pi_{0}$ (popular baseline). All results are summarized in Tab.3. In the Temporal Dimension, as the disturbance frequency decreases from Constant (1:0) to Sparse (1:3), all models exhibit improved R-Scores. Single-frame models ($\pi_{0}$, SpatialVLA and CogACT) tend to produce out-of-distribution actions under high-frequency disturbances, leading to task failure. RoboVLMs and TraceVLA, despite being multi-frame, heavily rely on accurate historical information and often default to inaction or repeated probing when disturbed. In contrast, CronusVLA demonstrates superior robustness due to effective multi-frame modeling, achieving strong resistance under Constant (1:0) and near immunity under Sparse (1:3) with 96.2 R-Score. In the Spatial Dimension, SpatialVLA is highly sensitive to local disturbances due to its reliance on zero-shot depth estimation, while CogACT and RoboVLMs are more affected by global disturbances. CronusVLA consistently outperforms others under these two types, exhibiting strong robustness. Notably, although SpatialVLA outperforms RoboVLMs on the original SimplerEnv benchmark, it underperforms on SimplerEnv-OR, with lower SR and R-Score, which highlights RoboVLMs’ robustness advantage as a multi-frame model. Overall, SimplerEnv-OR provides quantitative evidence of CronusVLA’s and other VLAs’ robustness under observational disturbances.
4.2 Real-world Experimental Results
@@
Evaluation setup within Franka platform. As shown in Fig.4, we evaluate our method on several real-world tasks with the Franka Research 3 Robot. We utilize a third-person camera for visual input. Three task suites are designed: (1) Simple Tasks, involves the picking and placing objects, opening and closing the drawer, and stacking cubes or cups; (2) Long-horizon Tasks, requires coordinated multi-step manipulation and includes putting multiple objects, placing objects into a drawer and pressing buttons in a specific order; and (3) Generalization and Observational Robustness tasks, evaluating performance on unseen objects, novel instructions, camera occlusions, distractor objects and so on. We manually collect 50 episodes for each task, and successful rollouts among 25 trials are reported. We implement DP3 (Ze et al. 2024), and OpenVLA (Kim et al. 2024) finetuned on these demonstrations, our CronusVLA 7B is finetuned from the post-trained weight. More details are in Appendix.
@@
Experimental results within Franka platform. CronusVLA outperforms other models across almost all tasks. For Simple Tasks, all three policies perform well when handling simple objects; however, for tasks requiring more precise manipulation, such as stacking one block (or cup) on another, CronusVLA shows better in-domain performance. For Long-horizon Tasks, CronusVLA exhibits stronger long-horizon learning capabilities, achieving consistently better performance than DP3 and OpenVLA under limited expert demonstrations. Notably, in press buttons in order, OpenVLA tends to press the same button multiple times, indicating state confusion in long-horizon tasks due to the absence of multi-frame information, while CronusVLA has inherent temporal awareness for button pressing. Generalization and Observational Robustness tasks are illustrated in Fig.4 (c) and (d). In all distracted situations, our model shows the best performance. DP3 is sensitive to distractions and human interference. In the Camera Occlusions, DP3 and OpenVLA are adversely affected by frequent visual dropouts due to their reliance on precise per-step observations, while our multi-frame modeling effectively withstands such disturbances.